A can do a piece of work in 20 days which B can do it in 12 days. B worked at it for 9 days. A can finish remaining work in:
a)3
b)5
c)7
d)11

Answer by mananth(16949)

B does the job in 12 days
so he does 1/12 of the job in 1 day.
so he does 9/12 of the job in 9 days = 3/4 of the job.
balance job will be 1/4
...
A does the job in 20 days
so he does 1/20 of the job in 1 day.
1/4 of the job he will take
..
(1/4)/(1/20)
1/4 *20 =5 days
m.ananth@hotmail.ca